Problem 15.47 4.0 kg of water at 39 OC is mixed with 4.0 kg of water at 52 °C in a well-insulated container Part A Estimate the net change in entropy of the system Express your answer to two significant figures and include the appropriate units. AS: Value Units Submit My Answers Give Up Provide eedback Continue
Answer #1

Heat capacity = Cp = 4184 J/kg.K

The change in entropy is given by:

\Delta S = C_pln{\frac{T_f}{T_i}}

and the final temperature can be calculated from the following formula:

T_f = \frac{m_1T_1+m_2T_2}{m_1+m_2}

T_1 = 273+39 = 312 K

T_2 = 273+52 = 325 K

Tf = (T_1+T_2)/2 = 318.5 K = 45.5 deg C

Hence net change in entropy = 4184*(ln(318.5/312)+ln(318.5/325)) =1.743 J/kg.K